OpenStudy (anonymous):

Someone please help! I don't get it. Suppose that x and y vary inversely and that x = 2 when y = 8. Write the function that models the inverse variation.

OpenStudy (the_fizicx99):

Y = k/x 8 = k/2 2(8) = 2(k/2) 16 = k 8 = 16/2 8 = 8

OpenStudy (anonymous):

thank you so much!!

OpenStudy (the_fizicx99):

The formula for inverse variation Y = k/x